OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation (x + 3)y = 14? Choose all answers that are correct. A. (−5, −7) B. (−1, 7) C. (10, −2) D. (11, 1)

OpenStudy (alexandervonhumboldt2):

ubstutute (x, y) coordinates to equation:

OpenStudy (alexandervonhumboldt2):

like (-5+3)*(-7)=14 and iwth each point

OpenStudy (anonymous):

x y 1 5 2 10 3 15 4 20 5 25 The table above represents what type of function? linear and nonlinear linear neither linear nor nonlinear nonlinear

OpenStudy (alexandervonhumboldt2):

5/1=5 10/2=5 15/3=5 so as the rate is same it is linear

OpenStudy (alexandervonhumboldt2):

also @poookieboo you can tag people so they will come to your question by typin @ and then start typing first letter of name

OpenStudy (alexandervonhumboldt2):

ans aslso you can medal person in one question post only once

OpenStudy (anonymous):

ok thanks

OpenStudy (alexandervonhumboldt2):

also by rules try posting one question per post

OpenStudy (anonymous):

ok srry